OpenStudy (anonymous):

Find the points of intersection of X^4-2x^2+1 and 1-X^2

OpenStudy (anonymous):

are these y=...

OpenStudy (dumbcow):

most likely in that case, set them equal to each other

OpenStudy (anonymous):

WHOAH! That graph is so weird!!!

OpenStudy (anonymous):

yes y =

OpenStudy (anonymous):

the parabola has a dent!

OpenStudy (anonymous):

i know lol

OpenStudy (dumbcow):

x^4-2x^2+1 = 1-x^2 x^4 -x^2 = 0 x^2(x^2 -1) = 0 x^2 = 0 --> x =0 x^2 - 1 =0 --> x = +-1

OpenStudy (anonymous):

my calc says 1,0 and -1,0

OpenStudy (anonymous):

thanks guys